OpenStudy (anonymous):

A mixture placed in an Erlenmeyer flask comprises of 6ml of silica gel and 40ml of solvent containing, in solution, 100mg of a non-volatile compound. After stirring, the mixture was left to stand before a 10ml aliquot of the solution was extracted and evaporated to dryness. The residue weighed 12mg. Calculate the adsorption coefficient K=C(stationary phase)/C (mobile phase) of the compound in this experiment.

OpenStudy (aaronq):

do you understand what's going on?

OpenStudy (anonymous):

Not really for this question. For the other chromatography question I can solve them using the equations given and relationships between them, but this one I have no clue.

OpenStudy (aaronq):

so you had 100 mg of a compound, 40 mL of solvent and 16 mL of gel. now, you mixed it and sampled a 10 mL portion of the solvent and found 12 mg of the compound. 10 mL is i quarter of the 40 mL total, so 12 mg*4= 48 mg of the solute is in the solvent, (100-48mg)=52 mg is adsorbed to the gel. can you finish it now?

OpenStudy (anonymous):

okay say 52mg is adsorbed into the gel what is the relationship between the math you did and the equation shown above?

OpenStudy (aaronq):

adsorption coefficient, K=C(stationary phase)/C (mobile phase) C = concentration stationary phase = gel mobile phase = solvent

OpenStudy (anonymous):

concentration = mol/vol i cant get mol

OpenStudy (aaronq):

concentration can be anything, you obviously can't use moles, use mg/mL